OpenStudy (anonymous):

how to convert 22.5 radians into pi form?

OpenStudy (tkhunny):

Ummm... a "radian" isn't really a thing. It's just a number. 22.5/pi = 7.16197243913529 Thus, 22.5 = 7.162pi (give or take)

OpenStudy (anonymous):

The options given are 13pi/8, pi/8, 9pi/8 or 5pi/8. The original question was find the supplement of the radian measure 7pi/8. Express answer in terms of pi.

OpenStudy (tkhunny):

It is SO much easier to ask the right question. Supplemental Angles add to \(\pi\). The supplement to 7\(\pi\)/8 is \(\pi\)/8. There is no point to converting to angle measure. What you should have asked, after pointlessly converting to degrees, is what is the Radian equivalent to 22.5º. That would be \(\pi\)/8.

OpenStudy (anonymous):

thank you for your help!